For what value of x will the rectangle at the right be a square?
Length = 2x+1 Width = x+2

Answer by ankor@dixie-net.com(22740)

For what value of x will the rectangle at the right be a square?
Length = 2x+1 Width = x+2
:
A square is Length = width, therefore
2x + 1 = x + 2
2x - x = 2 - 1
x = 1